OpenStudy (shelby1290):

Without graphing, determine whether each system has one solution, no solution or infinitely many solutions. Use slope and y-intercept. a) 3x + y = 4 6x + 2y = 8 b) x + 2y - 7 = 0 3x + 6y - 14 = 0

OpenStudy (anonymous):

a) y=3x+4 and y=3x+4 b) y=x-3.5 and y=2x- 7/3

OpenStudy (shelby1290):

@DerpyTurtle did you use slope and y-intercept to get those answers ^^^ and is it one solution, no solution or infinite?

OpenStudy (shelby1290):

please show me your work

OpenStudy (anonymous):

yes and a) no solution bc they r the same line b) i think one solution